OpenStudy (anonymous):

Find dy/dx of ln(x*y^2)=24 A. -2x/y B. -y/2x C. -y/2 D. -e^(-.5)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

could you explain how you did it? I tried, and I keep on getting the wrong answer.

OpenStudy (anonymous):

Solve the function for y, and get y = e^24/sqrt(x*e^24) Differentiate in respect to y

OpenStudy (anonymous):

oh, nvm. I accidentally messed up a multiplication.

OpenStudy (anonymous):

I get B

OpenStudy (anonymous):

\[dy/dx = (1/(x*y^2))(y^2+2yx(dy/dx)) =0\] \[dy/dx = -1/x*(y/2) = -y/(2x)\ = -2x/y] so yeah its A my bad

OpenStudy (anonymous):

\[dy/dx = -1/x*y/2 = -2x/y\] A